# UniBias: Unveiling and Mitigating LLM Bias through Internal Attention and FFN Manipulation

### v1

@@ 4.3 Alleviating Prompt Brittleness
Figure 1 presents Llama-2 7b’s performance both with and without UniBias. Without UniBias, the standard ICL’s performance varies significantly, ranging from 8% to 26%, demonstrating its instability. After applying UniBias, the accuracy stabilizes, with variations consistently less than 4% under perturbations of various design settings. This evidence verifies that UniBias effectively reduces prompt brittleness and enhances robustness.
@@
4.4 Biased LLM Components Analysis
In response to RQ3, we present the frequency counts of identified biased attention heads (AHs) and FFNs under repeated experiments in Figure 6. A large frequency count for an LLM component indicates a higher repeat of being identified as biased in the corresponding dataset. The first subfigure displays the biased components for various example selections, revealing several commonly biased LLM components across different prompts within a single dataset. The second subfigure highlights the common biased components across different datasets (ARC and MMLU) for the reasoning task, indicating that different datasets with similar tasks could share common biased LLM components. The third subfigure demonstrates the presence of common biased LLM components across different tasks. Experimental results suggest an interesting future direction: we may identify global biased components that would mitigate bias across multiple tasks and diverse prompt design settings.
4.5 Ablations
@@
Additionally, we further conduct experiments to investigate the impact of support set size (RQ 5), which is detailed in Appendix C.
@@ 5 Related Work
Bias in LLMs:   It is well recognized that LLMs are unstable under various ICL design settings, and this instability arises from biases in LLMs toward predicting certain answers [29, 14]. To understand these biases, existing studies have identified various bias factors, including recency bias, majority label bias, common token bias [29], and domain label bias [7] in classification tasks. More recently, selection bias, which consistently favors specific options in multiple-choice questions, has also been identified [30, 25]. To address these biases, several calibration methods have been proposed, including contextual calibration [29], domain-context calibration [7], and prototypical calibration [10]. However, these identified bias factors and calibration methods are derived from external observations or adjustments of LLM outputs, leaving the underlying mechanisms within LLMs that cause such biases poorly understood.
Mechanistic Interpretability:   Mechanistic interpretability [6, 23] aims to explain the internal processes in language models, facilitating the interpretation of the contributions of individual model components to the final prediction. Our work builds on the understanding of the residual stream [6], the logit lens [16], and the interpretation of LLM components in the vocabulary space [4, 8].
@@ A.1 Datasets
AGNews [28] 4 7600
TREC [22] 6 500
@@
ARC-Challenge [2] 4 1170
MMLU [12] 4 2000
COPA [19] 2 100
@@ A.2 Implementation Details
Experiments on internal mechanisms of biased factors: All experiments are conducted on Llama-2 7b model. For the vanilla label bias experiment, we projecting all FFN value vectors into the vocabulary space and sum the label logits for all FFN vectors whose label logits rank within the top $10$ over the vocabulary to calculate uncontextual accumulated FFN logits. We change different set of label words in prompt to derive the label prediction frequency of different label pairs. For the recency bias experiment, based on findings in [24], instead of the summed attention weights over the whole example, we adopt the sum of attention weights on label words of the example, e.g. "Answer: positive" as the effective attention weight on each example. For the selection bias experiment, we use zeroshot ARC dataset prompts in Table 6, and we use 12 samples for each class. The attention weight is also summed on label words instead of the whole option.
@@
Unibias:   In our method, all threshold values are determined through grid searching as described in the methodology section. Specifically, we use 20 samples per class as the support set for grid searching in all experiments. For each repetition of the experiment, the support set is randomly selected based on different random seeds. Additionally, to manipulate biased FFN vectors and attention heads, we create masks for the attention heads of all attention layers and adjust the FFN coefficient values and attention head masks using the hook operation. Additionally, we conduct the experiment on four A5000 GPUs. We will release our code upon acceptance to facilitate easy reproduction.
@@ Appendix B Limitation and Future Work
Appendix C Impact of support set size
@@
Appendix D Prompt Templates
The prompt templates used in this work are provided below. We generate few-shot ICL templates follow the template styles in [10, 7], as illustrated in Table 4.

### v2

@@ 4.3 Alleviating Prompt Brittleness
Figure 1 presents Llama-2 7b’s performance both with and without UniBias. Without UniBias, the standard ICL’s performance varies significantly, ranging from 8% to 26%, demonstrating its instability. After applying UniBias, the accuracy remains consistently high and stable, with variations consistently less than 4% under perturbations of various design settings. We provide further theoretical analysis on why UniBias can mitigate prompt brittleness and address various bias factors in Appendix G.
Table 2: Experiments on eliminating common biased components. Attention heads that are frequently identified as biased are removed from the original Llama-2 7b model.
SST2 MMLU COPA RTE MR Trec Avg.
ICL 87.226.03 41.732.25 67.602.30 66.217.30 89.371.83 72.9212.42 70.84
Unibias 94.540.62 44.830.24 69.002.74 67.656.44 92.190.37 80.803.17 74.84
Eliminating Common Biased Components 94.32 0.60 44.201.14 68.002.87 67.374.60 92.430.09 77.604.75 73.98
4.4 Biased Components Analysis and Common Biased Components Elimination
In response to RQ3, we present the frequency counts of identified biased attention heads (AHs) and FFNs under repeated experiments in Figure 6. A large frequency count for an LLM component indicates a higher repeat of being identified as biased in the corresponding dataset. The first subfigure displays the biased components for various example selections, revealing several commonly biased LLM components across different prompts within a single dataset. The second subfigure highlights the common biased components across different datasets (ARC and MMLU) for the reasoning task, indicating that different datasets with similar tasks could share common biased LLM components. The third subfigure demonstrates the presence of common biased LLM components across different tasks.
Experimental results suggest an interesting future direction: we may identify global biased components that would mitigate bias across multiple tasks and diverse prompt design settings. We conduct an preliminary experiment to explore the potential of eliminating common biased components. Specifically, we eliminate attention heads that are frequently identified as biased and apply this setting to diverse tasks, rather than handling each task individually. Experimental results in Table 2 demonstrate that although not as effective as our full Unibias method, eliminating common biased components outperforms the vanilla ICL by a large margin. Experiment details are in Appendix 5.
4.5 Ablations
@@
Additionally, we further conduct experiments to investigate the impact of support set size (RQ 5), which is detailed in Appendix E.
@@ 5 Related Work
Bias in LLMs:   It is well recognized that LLMs are unstable under various ICL design settings, and this instability arises from biases in LLMs toward predicting certain answers (Zhao et al., 2021; Lu et al., 2022). To understand these biases, existing studies have identified various bias factors, including recency bias, majority label bias, common token bias (Zhao et al., 2021), and domain label bias (Fei et al., 2023) in classification tasks. More recently, selection bias, which consistently favors specific options in multiple-choice questions, has also been identified (Zheng et al., 2023; Wang et al., 2023b). To address these biases, several calibration methods have been proposed, including contextual calibration (Zhao et al., 2021), domain-context calibration (Fei et al., 2023), and prototypical calibration (Han et al., 2023). However, these identified bias factors and calibration methods are derived from external observations or adjustments of LLM outputs, leaving the underlying mechanisms within LLMs that cause such biases poorly understood.
Prompt Brittleness:   Regarding prompt brittleness, it is demonstrated in the literature that this instability of prompt arises from LLMs’ inherent bias towards predicting certain answers (Zhao et al., 2021). Therefore, current research efforts address the prompt brittleness by mitigating LLMs’ bias towards labels (Fei et al., 2023; Han et al., 2023; Zhao et al., 2021).
Mechanistic Interpretability:   Mechanistic interpretability (Elhage et al., 2021; Wang et al., 2022) aims to explain the internal processes in language models, facilitating the interpretation of the contributions of individual model components to the final prediction. Our work builds on the understanding of the residual stream (Elhage et al., 2021), the logit lens (Nostalgebraist, 2020), and the interpretation of LLM components in the vocabulary space (Dar et al., 2023; Geva et al., 2021).
@@ 6 Conclusion
Acknowledgments and Disclosure of Funding
The authors would like to thank Edmond Lo, Lihui Chen, Xiyu Zhang, and the anonymous reviewers for their constructive comments and suggestions. The research was conducted at the Future Resilient Systems at the Singapore-ETH Centre, which was established collaboratively between ETH Zurich and the National Research Foundation Singapore. This research is supported by the National Research Foundation Singapore (NRF) under its Campus for Research Excellence and Technological Enterprise (CREATE) programme.
@@ A.1 Datasets
AGNews [Zhang et al., 2015] 4 7600
TREC [Voorhees and Tice, 2000] 6 500
@@
ARC-Challenge [Clark et al., 2018] 4 1170
MMLU [Hendrycks et al., 2020] 4 2000
COPA [Roemmele et al., 2011] 2 100
@@ A.2 Implementation Details
Experiments on internal mechanisms of biased factors: All experiments are conducted on Llama-2 7b model. For the vanilla label bias experiment, we projecting all FFN value vectors into the vocabulary space and sum the label logits for all FFN vectors whose label logits rank within the top $10$ over the vocabulary to calculate uncontextual accumulated FFN logits. We change different set of label words in prompt to derive the label prediction frequency of different label pairs. For the recency bias experiment, based on findings in [Wang et al., 2023a], instead of the summed attention weights over the whole example, we adopt the sum of attention weights on label words of the example, e.g. "Answer: positive" as the effective attention weight on each example. For the selection bias experiment, we use zeroshot ARC dataset prompts in Table 8, and we use 12 samples for each class. The attention weight is also summed on label words instead of the whole option.
@@
Unibias:   In our method, all threshold values are determined through grid searching as described in the methodology section. Specifically, we use 20 samples per class as the support set for grid searching in all experiments. For each repetition of the experiment, the support set is randomly selected based on different random seeds. Additionally, to manipulate biased FFN vectors and attention heads, we create masks for the attention heads of all attention layers and adjust the FFN coefficient values and attention head masks using the hook operation. Additionally, we conduct the experiment on four A5000 GPUs.
@@ Appendix B Limitation and Future Work
Appendix C Evaluation on More LLMs
Figure 7 demonstrates the performance of various methods across multiple datasets when applied to GPT-J and GPT2-XL models. For both models, our UniBias method consistently outperforms the baseline methods including vanilla ICL, CC, DC and PC. Notably, the improvement on the GPT2-XL model is substantial, demonstrating over an over $20\%$ increase in accuracy on SST-2 dataset compared to vanilla ICL.
Figure 7: Performance comparison of our UniBias method against baseline methods using GPT-J and GPT2-XL models.
Appendix D Eliminating Common Biased Components
Table 5: List of common biased attention heads eliminated. Indexing Starts from 0.
($19,10$) $(19,14)$ $(16,29)$ $(19,21)$ $(25,21)$ $(16,11)$ $(18,31)$ $(18,1)$
We explore the potential of eliminating common biased components and apply it to diverse tasks, rather than addressing each task individually. We conduct additional experiments on multiple tasks to assess the effectiveness of directly elinimate these components. Experimental results in Table 2 indicate that although not as effective as our full Unibias method, it outperforms the vanilla ICL by a large margin. Notably, eliminating common biased components represents cost-free gain in performance, as it involves only the direct masking of biased components identified in our work and is applicable to diverse tasks. The attention heads that are masked are listed in Table 5.
@@ Appendix D Eliminating Common Biased Components
Figure 9: Performance of Unibias using unlabeled samples as support set. It is compared against standard ICL and the original Unibias.
Appendix E Impact of Support Set Size
@@
Appendix F Using Unlabeled Samples for Support Set
To Address the potential challenge in accessing labeled samples, we further explore the alternative of using unlabeled samples during grid search. In our method, labeled samples are used to ensure each class is represented proportionally in the grid search, without direct use of the specific label information. Therefore, for balanced datasets, it is equally effective to employ a slight larger pool of unlabeled samples.
Our experimental findings, illustrated in Figure 9 of the rebuttal PDF, indicate that approximately $40\times\#Classes$ unlabeled samples achieves performance comparable to that obtained with labeled samples.
Appendix G Additional Analysis
We further analyze why mitigating model’s bias towards labels can alleviate prompt brittleness in our method. Due to the inherent bias of LLMs, different prompts can lead to varying biases towards labels. For example, due to recency bias, placing a negative sentiment analysis sample at the end of a prompt can make LLMs tend to predict ’negative’, incorrectly classifying positive samples and thus degrading ICL performance. Various bias factors lead to different direction and extend of bias, resulting in different changes in ICL performance and leading to the prompt brittleness. In contrast, our UniBias method effectively mitigates various potential biases inherent in LLMs by addressing their root causes internally from LLMs. By doing so, it minimizes the introduction of bias towards labels regardless of the difference in prompts, leading to more stable and accurate ICL performance across different prompt configurations.
Additionally, our UniBias method seeks to address a broad range of factors that lead to LLM bias, extending beyond those discussed in Section 2. Given the significant variability in prompts, models, and data corpuses, numerous unanticipated bias factors may emerge. Our approach is designed to tackle these diverse bias factors comprehensively. This is feasible because biased behaviors observed externally in LLMs originate from their internal components—specifically, the feedforward neural network (FFN) vectors and attention heads, which house nearly all LLM parameters. By directly identifying and mitigating biases within these FFN vectors and attention heads, UniBias offers a foundational strategy to counteract various forms of bias.
Appendix H Prompt Templates
The prompt templates used in this work are provided below. We generate few-shot ICL templates follow the template styles in [Han et al., 2023, Fei et al., 2023], as illustrated in Table 6.